GIVEN THE FUNCTION

F(x)= x2-16 / x-4 , x ≠ 4

7 , x=4

ESTABLISH IF IT IS CONTINUOUS IN X=4
Answer #1

Bol, flac) = x²16 x-4 24 ] 2=4 (x-4) (2+4) x & 4 Sa+4 1274 f(x) = e-4 2-4 7 1x=4 lim flo) = lem f4+h) = lim hoo (4th 4th +